Calculate the average translational kinetic energy, ?k, for one mole of gas at 649 K. Translational kinetic energy is sometimes called average kinetic energy.
Answer #1

We have ,Total K.E = n ( 3/2 R T )

Where, n is number of moles of gas, T is temoerature of a gas & R is gas constant.

\therefore Total K.E = 1 mol \times ( 3 /2 ) \times 8.314 J mol -1 K -1\times 649 K

Total K.E = 8039.7 J     

We have relation , 1 mol Compound = 6.022 \times 10 23 molecules

\therefore Average translational K.E = 8039.7 J / ( 6.022 \times 10 23 ) 1 = 1.344 \times 10 -20 J

ANSWER : Average translational K.E =1.344 \times 10 -20 J